# Sample Compression Scheme Reductions

### v1

@@ Appendix A Discussion about Infinitized Sample Compression Schemes
For binary concept classes with VC dimension $\mathrm{VC}$, Moran and Yehudayoff [2016] demonstrated the existence of a constant-size sample compression scheme of size $2^{O(\mathrm{VC})}$. However, we show that when the number of samples is infinite, such a result is no longer possible. Additionally, we construct an infinitized compression scheme for classes with finite Littlestone dimension. However, we also show that this is not a necessary condition—there are cases where the Littlestone dimension is infinite, yet the class still admits an infinitized compression scheme.
@@ Appendix B A Sample Compression Scheme for Classes with Graph Dimension $1$
Consider a binary class $\mathcal{C}$ on $\mathcal{X}$ with $\mathrm{VC}(\mathcal{C})\leq 1$. Pick any $c_{0}\in\mathcal{C}$. Define a partial ordering $\leq$ on $\mathcal{X}$ as follows: for $x,y\in\mathcal{X}$, let $x\leq y$ if, for every $c\in\mathcal{C}$, $c(x)\neq c_{0}(x)$ implies $c(y)\neq c_{0}(y)$. It holds that $\leq$ is a tree ordering. Furthermore, for any $c\in\mathcal{C}$, $\{x:c(x)\neq c_{0}(x)\}$ is a linear ordering.
@@
We introduce some notation and an algorithm that will be used to prove Lemma B.1. Consider a class $\mathcal{C}$ for which $\mathrm{d}_{\mathrm{G}}(\mathcal{C})=1$. Pick any $c_{0}\in\mathcal{C}$. For $c\in\mathcal{C}$, define $g_{c}$ such that $g_{c}(x)=\mathbbm{1}[c(x)\neq c_{0}(x)]$. Note the $g_{c_{0}}(x)=0$ for all $x$. Let $\mathcal{C}^{\prime}=\{g_{c}:c\in\mathcal{C}\}$. $\mathcal{C}^{\prime}$ is a binary class with VC dimension $\leq 1$. Consider points $x_{1},\ldots,x_{n}$. We can construct the tree ordering from Lemma B.3 with respect to $g_{c_{0}}$ on these points, where for any $x\leq y$ $g_{c}(x)\neq g_{c_{0}(x)}$ implies $g_{c}(y)\neq g_{c_{0}}(y)$. Since $g_{c_{0}}(x)$ is zero for all $x$, this implies that whenever $g_{c}(x)=1$, it must also hold that $g_{c}(y)=1$. By Lemma B.3, we have that for every $c\in\mathcal{C}$, the set $\{x:g_{c}(x)=1\}$ is a linear ordering. Below, we’ll propose what one may consider a natural first attempt at a compression scheme for multiclass, based on the binary compression scheme from Ben-David [2015].
@@
Since the $\mathrm{d}_{\mathrm{G}}(\mathcal{C})\leq 1$, the graph must be acyclic (Otherwise, consider a cycle. Since the graph is bipartite, the length of the cycle is at least $4$. Picking two disjoint edges $e_{1}=((z_{t},i),(z_{t+1},j)),e_{2}$, we can G-shatter $z_{t}$ and $z_{t+1}$ with labels $i$ and $j$ using $e_{1}$, $e_{2}$, and the two neighbors of $e_{1}$, contradicting that $\mathrm{d}_{\mathrm{G}}(\mathcal{C})\leq 1$). Thus, we can consider the graph to be an undirected forest. Furthermore, for any two disjoint edges, both of them must be incident to a leaf. (Otherwise, if there are two disjoint edges $e_{1},e_{2}$ that are disjoint, and $e_{1}$ is not incident with a leaf, then we can shatter the two points as earlier using the labels corresponding to $e_{1}$ with $e_{1}$, and the two neighbors of $e_{1}$).
@@
$[0,1]^{\mathcal{X}}$, it holds that $\mathrm{d}_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$.
@@
Applying Theorem C.1 gives that $\mathrm{d}_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$, i.e., the graph dimension of $\mathcal{C}$ is at most $4d_{\mathrm{P}}$. By our assumption, there exists a compression scheme of size $f(\mathrm{d}_{\mathrm{G}})$, so this implies that there exists a compression scheme of size $f(4d_{\mathrm{P}})$, as desired. The well-known bound of David et al. [2016] states that there is a bounded sample compression scheme of size $O(\mathrm{d}_{\mathrm{G}}(\mathcal{C})2^{\mathrm{d}_{\mathrm{G}}(\mathcal{C}})$. Using the reduction from pseudo-dimension to graph dimension, where $f(x)=cx2^{x}$ for some constant $c>0$, we get that there is a compression scheme of size $O(d_{\mathrm{P}}2^{4d_{\mathrm{P}}})$.
@@
There exists a partial concept class $\mathcal{C}_{part}$, such that $\mathrm{VC}(\mathcal{C}_{part})=1$, but there is no bounded-size sample compression scheme for $\mathcal{C}_{part}$.
@@
Suppose the graph has a cycle. For each edge in the cycle (if we traverse the cycle), the label of some $x_{i}$ will get flipped. Let $i_{1},\ldots,i_{k}$ be the sequence of indices that are flipped when we traverse the cycle, Let $i_{\ell 1}=i_{\ell_{2}}=a$, with $\ell_{1}\neq\ell_{2}$, $(\ell_{2}-\ell_{1}+k)\mod k$ minimal. Since the difference is minimal, traversing indices from $\ell_{1}+1$ to $\ell_{2}-1$ (modulo $k$) will traverse through distinct elements. Let one of these elements be $b$. Since this is a cycle, each element that is flipped needs to be flipped at least twice, so there exists an element traversing from $\ell_{2}+1\ldots\ell_{1}-1$ that is also equal to $b$. This would imply that a subsequence that is equal to $a,b,a,b$, so points $a$ and $b$ can be shattered by $\mathcal{C}_{part}$, contradicting that $\mathrm{VC}(\mathcal{\mathcal{C}}_{part})=1$.

### v2

@@ Appendix A Discussion about Infinitized Sample Compression Schemes
For binary concept classes with VC dimension $d_{\mathrm{VC}}$, Moran and Yehudayoff [2016] demonstrated the existence of a constant-size sample compression scheme of size $2^{O(d_{\mathrm{VC}})}$. However, we show that when the number of samples is infinite, such a result is no longer possible. Additionally, we construct an infinitized compression scheme for classes with finite Littlestone dimension. However, we also show that this is not a necessary condition—there are cases where the Littlestone dimension is infinite, yet the class still admits an infinitized compression scheme.
@@ Appendix B A Sample Compression Scheme for Classes with Graph Dimension $1$
Consider a binary class $\mathcal{C}$ on $\mathcal{X}$ with $d_{\mathrm{VC}}(\mathcal{C})\leq 1$. Pick any $c_{0}\in\mathcal{C}$. Define a partial ordering $\leq$ on $\mathcal{X}$ as follows: for $x,y\in\mathcal{X}$, let $x\leq y$ if, for every $c\in\mathcal{C}$, $c(x)\neq c_{0}(x)$ implies $c(y)\neq c_{0}(y)$. It holds that $\leq$ is a tree ordering. Furthermore, for any $c\in\mathcal{C}$, $\{x:c(x)\neq c_{0}(x)\}$ is a linear ordering.
@@
We introduce some notation and an algorithm that will be used to prove Lemma B.1. Consider a class $\mathcal{C}$ for which $d_{\mathrm{G}}(\mathcal{C})=1$. Pick any $c_{0}\in\mathcal{C}$. For $c\in\mathcal{C}$, define $g_{c}$ such that $g_{c}(x)=\mathbbm{1}[c(x)\neq c_{0}(x)]$. Note the $g_{c_{0}}(x)=0$ for all $x$. Let $\mathcal{C}^{\prime}=\{g_{c}:c\in\mathcal{C}\}$. $\mathcal{C}^{\prime}$ is a binary class with VC dimension $\leq 1$. Consider points $x_{1},\ldots,x_{n}$. We can construct the tree ordering from Lemma B.3 with respect to $g_{c_{0}}$ on these points, where for any $x\leq y$ $g_{c}(x)\neq g_{c_{0}(x)}$ implies $g_{c}(y)\neq g_{c_{0}}(y)$. Since $g_{c_{0}}(x)$ is zero for all $x$, this implies that whenever $g_{c}(x)=1$, it must also hold that $g_{c}(y)=1$. By Lemma B.3, we have that for every $c\in\mathcal{C}$, the set $\{x:g_{c}(x)=1\}$ is a linear ordering. Below, we’ll propose what one may consider a natural first attempt at a compression scheme for multiclass, based on the binary compression scheme from Ben-David [2015].
@@
Since the $d_{\mathrm{G}}(\mathcal{C})\leq 1$, the graph must be acyclic (Otherwise, consider a cycle. Since the graph is bipartite, the length of the cycle is at least $4$. Picking two disjoint edges $e_{1}=((z_{t},i),(z_{t+1},j)),e_{2}$, we can G-shatter $z_{t}$ and $z_{t+1}$ with labels $i$ and $j$ using $e_{1}$, $e_{2}$, and the two neighbors of $e_{1}$, contradicting that $d_{\mathrm{G}}(\mathcal{C})\leq 1$). Thus, we can consider the graph to be an undirected forest. Furthermore, for any two disjoint edges, both of them must be incident to a leaf. (Otherwise, if there are two disjoint edges $e_{1},e_{2}$ that are disjoint, and $e_{1}$ is not incident with a leaf, then we can shatter the two points as earlier using the labels corresponding to $e_{1}$ with $e_{1}$, and the two neighbors of $e_{1}$).
@@
$[0,1]^{\mathcal{X}}$, it holds that $d_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$.
@@
Applying Theorem C.1 gives that $d_{\mathrm{G}}(\mathcal{C})\leq 4d_{\mathrm{P}}(\mathcal{C})$, i.e., the graph dimension of $\mathcal{C}$ is at most $4d_{\mathrm{P}}$. By our assumption, there exists a compression scheme of size $f(d_{\mathrm{G}})$, so this implies that there exists a compression scheme of size $f(4d_{\mathrm{P}})$, as desired. The well-known bound of David et al. [2016] states that there is a bounded sample compression scheme of size $O(d_{\mathrm{G}}(\mathcal{C})2^{d_{\mathrm{G}}(\mathcal{C}})$. Using the reduction from pseudo-dimension to graph dimension, where $f(x)=cx2^{x}$ for some constant $c>0$, we get that there is a compression scheme of size $O(d_{\mathrm{P}}2^{4d_{\mathrm{P}}})$.
@@
There exists a partial concept class $\mathcal{C}_{part}$, such that $d_{\mathrm{VC}}(\mathcal{C}_{part})=1$, but there is no bounded-size sample compression scheme for $\mathcal{C}_{part}$.
@@
Suppose the graph has a cycle. For each edge in the cycle (if we traverse the cycle), the label of some $x_{i}$ will get flipped. Let $i_{1},\ldots,i_{k}$ be the sequence of indices that are flipped when we traverse the cycle, Let $i_{\ell 1}=i_{\ell_{2}}=a$, with $\ell_{1}\neq\ell_{2}$, $(\ell_{2}-\ell_{1}+k)\mod k$ minimal. Since the difference is minimal, traversing indices from $\ell_{1}+1$ to $\ell_{2}-1$ (modulo $k$) will traverse through distinct elements. Let one of these elements be $b$. Since this is a cycle, each element that is flipped needs to be flipped at least twice, so there exists an element traversing from $\ell_{2}+1\ldots\ell_{1}-1$ that is also equal to $b$. This would imply that a subsequence that is equal to $a,b,a,b$, so points $a$ and $b$ can be shattered by $\mathcal{C}_{part}$, contradicting that $d_{\mathrm{VC}}(\mathcal{\mathcal{C}}_{part})=1$.